OpenStudy (anonymous):

solve each system of equations. 2y^2-3xy+6y+2x+4=0 2x-3y+4=0

OpenStudy (radar):

2x-3y+4=0, 2x=3y-4, x=(3/2)y-2 now substitute for in the 2nd degree equation. This gives you the following: \[2y ^{2}-(9/2)y ^{2}+6y+6y+3y-4+4=0\]\[(-5/2)y ^{2}+15y=0\] Multiply both sides by a -2 getting:\[5y ^{2}-30y=0\] Factor getting: 5y(y-6)=0 Solve for y getting: y=0 y=6

OpenStudy (radar):

Solving for x: 2x-0+4=0 2x=-4 x=-2 Other value 2x-18+4=0 2x=14 x=7

OpenStudy (radar):

Did you follow with understanding?

OpenStudy (anonymous):

yes!! thank you so much!!

OpenStudy (radar):

Good luck with your studies.